Sarah, age 47, is married with two children in high school. She estimates that her average annual earnings over the next 18 years will be $95,000. She estimates that 40% of her average annual earnings will be used to pay taxes, insurance premiums, and the cost of self-maintenance. The remainder will be used to support the family. Sarah wants to calculate the human life value and believes that a 5.5% discount rate is appropriate. Calculate Sarah's human life value.

Round your answer to the nearest dollar.
Answer #1

Amount used to support the family = 95000*(1-40%) = 57000 per annum

The number of years = 18

Discount rate = 5.5%

Human life value = Annual amount*(1-1/(1+r)^n)/r

= 57000*(1-1/1.055^18)/ 0.055

=$ 641026